# Mihai Dascalub, Traian Rebedeab,f

### v2

@@ Abstract
Mihai Masalaa,b, Denis C. Ilie-Ablachimb, Alexandru Dimaa,b, Dragos Corlatescub, Miruna Zavelcaa,c, Ovio Olarue, Simina Teriane, Andrei Teriane, Marius Leordeanub, Horia Velicud, Marius Popescuc, Mihai Dascalub, Traian Rebedeab,f aInstitute for Logic and Data Science, Bucharest, bNational University of Science and Technology POLITEHNICA Bucharest, Romania, cUniversity of Bucharest, Romania, dBRD Groupe Societe Generale, e Lucian Blaga University Sibiu, fNVIDIA
@@ 3.2 Instructions and Conversations
We collect a large and diverse dataset of conversations and instructions to build the instruct models. Since high-quality data in Romanian is very scarce, we resort to translating datasets such as Alpaca Taori et al. (2023), Dolly Conover et al. (2023), NoRobots Rajani et al. (2023), GPT-Instruct777www.github.com/teknium1/GPTeacher, Orca Mukherjee et al. (2023) and Camel Li et al. (2023a). This led to a total of around 650k translated instructions888www.systransoft.com/ last accessed 6th of March 2024, which were further used in our experiments.
@@ 4.1 Foundational Model
Training the foundational model is done iteratively using continual pre-training on the CulturaX dataset. Starting from the foundational Llama2 model, we trained the first model on 5% of the CulturaX, the following model on the next 5% (leading to the second model seeing 10% of the data), and the third model on another 10% of the CulturaX. Therefore, we have three variants of a Romanian foundational model, variants that used 5%, 10%, and 20% of the entire CulturaX dataset. Training is done over one epoch, with a fixed learning rate of $10^{-4}$.
@@ 5 Evaluation
Llama-2-7b-chat 36.87 37.06 33.81 56.02 45.36 4.88 44.09
RoLlama2-7b-Instruct 45.71 43.66 39.70 70.34 57.36 18.78 44.44
@@
Mistral-7B-v0.1 45.11 42.88 47.29 60.75 54.17 20.39 48.77
Mistral-7B-Instruct-v0.2 47.41 46.25 47.04 58.72 54.25 13.59 64.63
RoMistral-7b-Instruct 52.49 50.39 51.64 66.69 60.24 33.71 52.29
@@
Llama-3-8B 45.35 42.19 50.41 59.73 53.46 20.39 45.90
Llama-3-8B-Instruct 50.15 43.73 49.02 59.35 53.16 44.12 51.52
RoLlama3-8b-Instruct 50.61 44.66 52.19 67.58 57.65 30.20 51.39
@@ 5.3 Romanian Downstream Tasks
Llama-2-7b-chat 0.18 0.14 0.23 5 / 160
Llama-2-7b-chat (ro_prompted) 1.21 1.68 0.74 44 / 160
RoLlama2-7b-Instruct 3.70 4.74 2.66 160 / 160
Mistral-7B-Instruct-v0.2 0.87 1.15 0.59 25 / 160
Mistral-7B-Instruct-v0.2 (ro_prompted) 4.83 5.09 4.58 154 / 160
RoMistral-Instruct 4.91 5.67 4.16 160 / 160
Llama-3-8B-Instruct 4.94 5.41 4.48 136 / 160
Llama-3-8B-Instruct (ro_prompted) 5.92 6.36 5.49 158 / 160
RoLlama3-8b-Instruct 5.28 6.10 4.45 160 / 160
gemma-1.1-7b-it 4.89 5.54 4.25 158 / 160
gemma-1.1-7b-it (ro_prompted) 4.63 5.18 4.08 160 / 160
RoGemma-7b-Instruct 4.83 5.56 4.10 160 / 160
Table 2: Performance of Instruct models on Romanian MT-Bench. Bold as in Table 1.
@@
Llama-2-7b 88.82 64.30 98.43 87.22
RoLlama2-7b-Base 91.98 63.89 98.97 87.72
Llama-2-7b-chat 80.67 57.51 97.27 82.02
RoLlama2-7b-Instruct 96.27 66.42 98.83 87.28
Mistral-7B-Instruct-v0.2 97.00 62.08 98.83 87.32
RoMistral-7b-Instruct 97.36 67.55 98.80 88.28
Llama-3-8B-Instruct 95.71 59.04 98.53 86.19
RoLlama3-8b-Instruct 97.52 67.41 94.15 87.13
gemma-1.1-7b-it 87.76 60.17 83.87 85.61
RoGemma-7b-Instruct 97.87 65.71 98.43 87.17
@@
Llama-2-7b-chat 15.56 28.51 19.99 31.48
RoLlama2-7b-Instruct 27.39 10.32 27.59 40.13
Mistral-7B-Instruct-v0.2 18.58 34.04 26.19 39.88
RoMistral-7b-Instruct 27.93 13.22 28.72 40.86
Llama-3-8B-Instruct 19.02 30.98 28.02 40.28
RoLlama3-8b-Instruct 24.02 27.37 26.53 40.36
gemma-1.1-7b-it 17.91 27.82 25.48 36.11
RoGemma-7b-Instruct 27.91 23.08 27.99 39.51
@@
Llama-2-7b-chat 32.29 53.98 60.34 75.98
RoLlama2-7b-Instruct 44.52 64.75 54.96 70.20
Mistral-7B-Instruct-v0.2 27.71 50.43 65.46 79.73
RoMistral-7b-Instruct 43.66 63.70 55.04 72.31
Llama-3-8B-Instruct 39.58 58.77 67.65 82.77
RoLlama3-8b-Instruct 39.43 59.50 44.45 59.76
gemma-1.1-7b-it 42.18 62.43 60.34 77.40
RoGemma-7b-Instruct 17.75 28.11 52.02 68.43
@@
Llama-2-7b-chat 32.12 32.64 74.08 72.64
RoLlama2-7b-Instruct 65.50 67.79 84.44 84.76
Mistral-7B-Instruct-v0.2 61.08 62.83 84.92 85.44
RoMistral-7b-Instruct 77.43 78.43 87.25 87.79
Llama-3-8B-Instruct 72.35 72.82 83.49 84.06
RoLlama3-8b-Instruct 77.20 77.87 85.80 86.05
gemma-1.1-7b-it 50.03 48.72 83.43 83.64
RoGemma-7b-Instruct 73.96 75.16 86.45 86.31
@@ 5.4 RoCulturaBench
We provide reference answers for all prompts and evaluate the LLM-generated answers using GPT-4o as a judge. Compared to MT-Bench, RoCulturaBench contains only single-turn interactions (i.e., one instruction, one answer).
@@ 6.1 Academic Benchmarks
Table 1 introduces the academic benchmark results. In the first section, Llama2, and different versions of RoLlama2 foundational model are compared. Increasing the pretraining sequence length boosts the overall performance. Adding more pretraining data (up to 20% of our data) degrades the foundational model’s performance. We conjecture that this is due to the quality of the pre-training data used, as CulturaX contains data from different sources, including low-quality sources. Nevertheless, we found foundational models exposed to more training data responding better to instruction fine-tuning, consistently outperforming models trained on less data. For that reason, even if the model has lower performance than some of its counterparts, we perform instruction fine-tuning on RoLlama2-7b-Base that was trained on 20% of the data with a sequence length of 1024. Still, pre-training on Romanian texts increases the performance on four of the six benchmarks when compared to the Llama2 foundational model. There is a noticeable increase in the Winogrande and HellaSwag benchmarks, arguing for the model’s capability to understand Romanian and use it for reasoning. On GSM8k and MMLU, we observe a consistent drop in performance that could be explained by the lack of appropriate data in pre-training, thus making it easier for the model to forget mathematical reasoning (for GSM8k) and knowledge about US history or law. Fine-tuning with instructions further increases the overall score by around 7% over RoLlama2-base with an increase on all of the six benchmarks.
@@ 6.2 MT-Bench
Performance on MT-Bench is presented in Table 2. We noticed that even when using a Romanian prompt, most Llama2, Mistral, and Llama3 answers, while somewhat correct, are either entirely in English or have the first few words in Romanian, then reverting to English. We specifically ask the judge to penalize answers that are not in Romanian, but unfortunately, experiments showed that the GPT judge scarcely does this, assigning the same score to English answers irrespective of our request. For this reason, we decided to automatically detect the language of the answers111111www.pypi.org/project/langdetect/ and penalize the answers that are not written in Romanian with a score of 0. Upon manual inspection, we decided to manually assign the language using simple rules for a few specific prompts where automated methods fail. To alleviate the predisposition of models to answer in English even if the prompt is in Romanian, we added after each instruction an additional request to answer only in Romanian. As seen in the last column in Table 2, when specifically prompted to answer in Romanian (ro_prompted), the models tended to perform better, with Llama3 and Gemma being capable of answering mostly in the correct language. As expected, this is not an issue with RoLLMs, as they generate text in Romanian without being specifically told to do so.
@@
Llama-2-7b-chat (ro_prompted) 1.72 48 / 100
RoLlama2-7b-Instruct 3.43 100 / 100
Mistral-7B-Instruct-v0.2 (ro_prompted) 3.75 99 / 100
RoMistral-Instruct 3.17 100 / 100
Llama-3-8B-Instruct (ro_prompted) 4.61 100 / 100
RoLlama3-8b-Instruct 3.83 100 / 100
gemma-1.1-7b-it (ro_prompted) 3.22 100 / 100
RoGemma-7b-Instruct 3.47 100 / 100
GPT-3.5 5.55 100 / 100
Table 7: Performance of Instruct models on RoCulturaBench. Bold denotes as in Table 1.
Turning to the evaluation scores, we notice the gap in performance between the 1st and 2nd turns is larger for RoLLMs than for their counterparts. For RoLLMs, this is somewhat expected as the instruction fine-tuning dataset contains, for the most part, single-turn conversations. Adding more multi-turn conversations in the fine-tuning dataset should also close the performance gap on the second turn, thus leading to greater overall performance. We also emphasize that we performed instruction fine-tuning on base models, thus effectively losing all the valuable time and resources spent on alignment. For example, Llama3-Instruct has been fine-tuned and aligned with human preference on a huge dataset that also contains 10M human-annotated examples. To put things into perspective, we translated and used around 650k English instructions, out of which only 25k are human-created. We leave the investigation of alignment methods for future work. Furthermore, Llama3, the best-performing model on Romanian MT-Bench is a recently released model, and we are still investigating how to best train such a model, including continual instruction fine-tuning Chen and Lee (2024).
@@ 6.3 Romanian Downstream Tasks
In the case of question answering (see Table 5), the fine-tuning setup is different due to the nature of the training data as the XQuAD dataset consists of 1190 question-answer pairs professionally translated from the development set of SQuAD v1.1 Rajpurkar et al. (2016). Therefore, fine-tuning in this case is done on the English , while evaluation is done on the Romanian question-answer pairs. Naturally, non-RO LLMs tend to respond better to data in English. Also, we note that the foundational models perform better than instruct variants in this particular case of XQuAD. In zero/few-shot settings, RoLLMs perform better, with the exception of RoGemma-7b-Instruct, which has very poor performance on 3 and 5-shot settings.
@@ 6.4 RoCulturaBench
Results on RoCulturaBench are presented in Table 7. Taking an English foundational model and adapting it to a new language by performing instruction fine-tuning is not enough to properly understand the nuances and realities of a culture. Testament stands the competitive performance of RoLlama2: starting from an arguably weaker model but also pre-training it performs on par with newer and stronger models that have only been fine-tuned. We argue that, as expected, cultural information and social knowledge are primarily injected into the model in the pre-training phase, not in the fine-tuning phase, especially when fine-tuning exploits non-natively data (i.e., written in English and then translated). As in the case of MT-Bench, Llama3 outperforms other models by a noteworthy margin, with only GPT-3.5 being stronger. For more detailed results and comparisons, see Appendix Section A.2.
We argue that results obtained using the LLM-as-a-judge approach are not the be-all and end-all of evaluations, especially regarding non-English language and culture. We leave human evaluation and other approaches (e.g., training a language-specific judge) for future work.

### v4

# A Recipe to Train Powerful Romanian LLMs with English Instructions

@@ Abstract
A Recipe to Train Powerful Romanian LLMs with English Instructions
Mihai Masalaa,b,c, Denis C. Ilie-Ablachimb, Alexandru Dimaa,b, Dragos Corlatescub, Miruna Zavelcaa,d, Ovio Olaruf, Simina Terianf, Andrei Terianf, Marius Leordeanub,c, Horia Velicue, Marius Popescud, Mihai Dascalub, Traian Rebedeab,g aInstitute for Logic and Data Science, Bucharest, bNational University of Science and Technology POLITEHNICA Bucharest, Romania, cSimion Stoilow Institute of Mathematics of The Romanian Academy, Bucharest, dUniversity of Bucharest, Romania, eBRD Groupe Societe Generale, fLucian Blaga University Sibiu, gNVIDIA
@@ 3.2 Instructions and Conversations
We collect a large and diverse dataset of conversations and instructions to build the instruct models. Since high-quality data in Romanian is very scarce, we resort to translating777www.systransoft.com/ last accessed 6th of March 2024 datasets such as Alpaca Taori et al. (2023), Dolly Conover et al. (2023), NoRobots Rajani et al. (2023), GPT-Instruct888www.github.com/teknium1/GPTeacher, Orca Mukherjee et al. (2023), Camel Li et al. (2023a), and UltraChat Ding et al. (2023). Furthermore, we use Romanian data from OpenAssistant Köpf et al. (2023) and Aya Collection Singh et al. (2024). This led to a total of around 2.7M instructions and conversations in Romanian, which were further used in our experiments.
@@ 4.1 Foundational Model
Training the foundational model is done iteratively using continual pre-training on the CulturaX dataset. Starting from the foundational Llama2 model, we trained the first model on 5% of the CulturaX, the following model on the next 5% (leading to the second model seeing 10% of the data), and the third model on another 10% of the CulturaX. Therefore, we have three variants of a Romanian foundational model, variants that used 5%, 10%, and 20% of the entire CulturaX dataset. Training is done over one epoch, with a fixed learning rate of $1e^{-4}$.
@@ 5 Evaluation
Llama-2-7b-chat 36.84 37.03 33.81 55.87 45.36 4.90 44.09
RoLlama2-7b-Instruct 44.50 44.73 40.39 63.67 59.12 13.29 45.78
@@
Mistral-7B-v0.1 45.02 42.99 47.16 60.77 54.19 16.20 48.80
Mistral-7B-Instruct-v0.2 47.40 46.29 47.01 58.78 54.27 13.47 64.59
RoMistral-7b-Instruct 52.91 52.27 49.33 70.03 62.88 32.42 50.51
@@
Llama-3-8B 44.55 38.05 48.33 59.94 53.48 20.04 47.44
Llama-3-8B-Instruct 50.62 43.69 52.04 59.33 53.19 43.87 51.59
RoLlama3-8b-Instruct 52.21 47.95 53.50 66.06 59.72 40.16 45.90
Llama3.1
Llama-3.1-8B 37.29 33.25 36.35 58.80 42.65 3.59 49.03
Llama-3.1-8B-Instruct 49.87 42.86 53.73 59.71 56.82 35.56 50.54
RoLlama3.1-8b-Instruct 53.03 47.69 54.57 65.85 59.94 44.30 45.82
@@ 5.3 Romanian Downstream Tasks
Llama-2-7b-chat 0.59 0.95 0.23 21 / 160
Llama-2-7b-chat (ro_prompted) 1.08 1.44 0.73 45 / 160
RoLlama2-7b-Instruct 4.43 4.93 3.94 160 / 160
Mistral-7B-Instruct-v0.2 1.04 1.35 0.72 27 / 160
Mistral-7B-Instruct-v0.2 (ro_prompted) 5.03 5.05 5.00 154 / 160
RoMistral-7b-Instruct 5.29 5.86 4.73 160 / 160
Llama-3-8B-Instruct 4.54 4.65 4.43 126 / 160
Llama-3-8B-Instruct (ro_prompted) 5.96 6.16 5.76 158 / 160
RoLlama3-8b-Instruct 5.38 6.09 4.68 160 / 160
Llama-3.1-8B-Instruct 5.69 5.96 5.43 160 / 160
Llama-3.1-8B-Instruct (ro_prompted) 5.69 5.85 5.53 160 / 160
RoLlama3.1-8b-Instruct 5.42 5.95 4.89 160 / 160
gemma-1.1-7b-it 5.09 5.53 4.66 158 / 160
gemma-1.1-7b-it (ro_prompted) 4.83 5.11 4.55 160 / 160
RoGemma-7b-Instruct 5.24 5.55 4.94 160 / 160
Table 2: Romanian MT-Bench performance. Bold as in Table 1.
@@
Llama-2-7b 93.19 54.11 98.43 87.22
RoLlama2-7b-Base 83.25 61.04 98.97 87.72
Llama-2-7b-chat 87.78 52.82 97.27 82.02
RoLlama2-7b-Instruct 97.66 62.41 97.97 60.89
Mistral-7B-Instruct-v0.2 96.97 56.66 98.83 87.32
RoMistral-7b-Instruct 95.56 67.83 99.00 87.57
Llama-3-8B-Instruct 95.88 56.21 98.53 86.19
RoLlama3-8b-Instruct 95.58 61.20 96.46 87.26
Llama-3.1-8B-Instruct 95.74 59.49 98.57 82.41
RoLlama3.1-8b-Instruct 94.57 60.10 95.12 87.53
gemma-1.1-7b-it 87.54 51.49 83.87 85.61
RoGemma-7b-Instruct 86.96 56.72 98.80 85.81
@@
Llama-2-7b-chat 15.55 28.53 19.99 31.48
RoLlama2-7b-Instruct 27.14 19.40 27.63 39.75
Mistral-7B-Instruct-v0.2 18.60 33.99 26.19 39.88
RoMistral-7b-Instruct 28.28 6.10 27.70 40.36
Llama-3-8B-Instruct 18.89 30.98 28.02 40.28
RoLlama3-8b-Instruct 22.92 24.28 27.31 40.52
Llama-3.1-8B-Instruct 19.01 27.77 29.02 39.80
RoLlama3.1-8b-Instruct 21.88 23.99 28.27 40.44
gemma-1.1-7b-it 17.96 27.74 25.48 36.11
RoGemma-7b-Instruct 24.45 14.20 25.96 39.07
@@
Llama-2-7b-chat 32.35 54.00 60.34 75.98
RoLlama2-7b-Instruct 45.71 65.08 59.24 74.25
Mistral-7B-Instruct-v0.2 27.92 50.71 65.46 79.73
RoMistral-7b-Instruct 41.09 63.21 47.56 62.69
Llama-3-8B-Instruct 39.47 58.67 67.65 82.77
RoLlama3-8b-Instruct 18.89 31.79 50.84 65.18
Llama-3.1-8B-Instruct 44.96 64.45 69.50 84.31
RoLlama3.1-8b-Instruct 13.59 23.56 49.41 62.93
gemma-1.1-7b-it 42.10 62.30 60.34 77.40
RoGemma-7b-Instruct 26.03 41.58 46.72 60.79
@@
Llama-2-7b-chat 32.56 31.99 74.08 72.64
RoLlama2-7b-Instruct 59.69 57.16 84.66 85.07
Mistral-7B-Instruct-v0.2 62.62 60.86 84.92 85.44
RoMistral-7b-Instruct 78.47 77.24 87.28 87.88
Llama-3-8B-Instruct 73.04 72.36 83.49 84.06
RoLlama3-8b-Instruct 77.60 76.86 86.70 87.09
Llama-3.1-8B-Instruct 72.11 71.64 84.59 84.96
RoLlama3.1-8b-Instruct 75.89 76.00 86.86 87.05
gemma-1.1-7b-it 49.10 50.23 83.43 83.64
RoGemma-7b-Instruct 73.23 71.58 88.42 88.45
@@ 5.4 RoCulturaBench
We provide reference answers for all prompts and evaluate the LLM-generated answers using GPT-4o111111gpt-4o-2024-08-06 as a judge. Compared to MT-Bench, RoCulturaBench contains only single-turn interactions (i.e., one instruction, one answer).
@@ 6.1 Academic Benchmarks
Table 1 introduces the academic benchmark results. In the first section, Llama2, and different versions of RoLlama2 foundational model are compared. Increasing the pretraining sequence length boosts the overall performance while adding more pretraining data (up to 20% of our data) slighty degrades the foundational model’s performance. We conjecture that this is due to the quality of the pre-training data used, as CulturaX contains data from different sources, including low-quality sources. Nevertheless, we found foundational models exposed to more training data responding better to instruction fine-tuning, consistently outperforming models trained on less data. For that reason, even if the model has lower performance than some of its counterparts, we perform instruction fine-tuning on RoLlama2-7b-Base that was trained on 20% of the data with a sequence length of 1024. Still, pre-training on Romanian texts increases the performance on four of the six benchmarks when compared to the Llama2 foundational model. There is a noticeable increase in the Winogrande and HellaSwag benchmarks, arguing for the model’s capability to understand Romanian and use it for reasoning. On GSM8k and MMLU, we observe a consistent drop in performance that could be explained by the lack of appropriate data in pre-training, thus making it easier for the model to forget mathematical reasoning (for GSM8k) and knowledge about US history or law. Fine-tuning with instructions further increases the overall score by around 6% over RoLlama2-base with an increase on all of the six benchmarks.
@@ 6.2 MT-Bench
Performance on MT-Bench is presented in Table 2. We noticed that even when using a Romanian prompt, most Llama2, Mistral, and Llama3 answers, while somewhat correct, are either entirely in English or have the first few words in Romanian, then reverting to English. We specifically ask the judge to penalize answers that are not in Romanian, but unfortunately, experiments showed that the GPT judge scarcely does this, assigning the same score to English answers irrespective of our request. For this reason, we decided to automatically detect the language of the answers121212www.pypi.org/project/langdetect/ and penalize the answers that are not written in Romanian with a score of 0. Upon manual inspection, we decided to manually assign the language using simple rules for a few specific prompts where automated methods fail. To alleviate the predisposition of models to answer in English even if the prompt is in Romanian, we added after each instruction an additional request to answer only in Romanian. As seen in the last column in Table 2, when specifically prompted to answer in Romanian (ro_prompted), the models tended to perform better, with Llama3 and Gemma being capable of answering mostly in the correct language. As expected, this is not an issue with RoLLMs, as they generate text in Romanian without being specifically told to do so.
@@
Llama-2-7b-chat (ro_prompted) 1.21 33 / 100
RoLlama2-7b-Instruct 4.08 100 / 100
Mistral-7B-Instruct-v0.2 (ro_prompted) 3.68 97 / 100
RoMistral-7b-Instruct 3.99 100 / 100
Llama-3-8B-Instruct (ro_prompted) 4.62 100 / 100
RoLlama3-8b-Instruct 3.81 100 / 100
Llama-3.1-8B-Instruct (ro_prompted) 3.54 100 / 100
RoLlama3.1-8b-Instruct 3.55 100 / 100
gemma-1.1-7b-it (ro_prompted) 3.38 100 / 100
RoGemma-7b-Instruct 3.51 100 / 100
GPT-3.5 6.09 100 / 100
Table 7: RoCulturaBench performance. Bold denotes as in Table 1.
Turning to the evaluation scores, we notice the gap in performance between the 1st and 2nd turns is larger for RoLLMs than for their counterparts. For RoLLMs, this is somewhat expected as the instruction fine-tuning dataset contains, for the most part, single-turn conversations. Adding more multi-turn conversations in the fine-tuning dataset should also close the performance gap on the second turn, thus leading to greater overall performance.
We also emphasize that we performed instruction fine-tuning on base models except for the newer Llama3 and Llama3.1 models, for which we lost valuable time and resources spent on alignment. For the Llama3 model family, we found better results by performing direct instruction fine-tuning on the Instruct variant, a model fine-tuned and aligned with human preferences on a huge dataset that also contains 10M human-annotated examples [^13] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, last accessed on 1st of October 2024. To put things into perspective, we translated and used around 2.7M English instructions, out of which only 25k were human-created. As such, we are still investigating how to best train such a model, including continual instruction fine-tuning Chen and Lee (2024).
@@ 6.3 Romanian Downstream Tasks
In the case of question answering (see Table 5), the fine-tuning setup is different due to the nature of the training data as the XQuAD dataset consists of 1190 question-answer pairs professionally translated from the development set of SQuAD v1.1 Rajpurkar et al. (2016). Therefore, fine-tuning in this case is done on the English dataset, while evaluation is done on the Romanian question-answer pairs. Naturally, non-RO LLMs tend to respond better to data in English. Also, we note that the foundational models perform better than instruct variants in this particular case of XQuAD. In zero/few-shot settings, RoLLMs perform better, with the exception of RoGemma-7b-Instruct and RoLlama3-8b-Instruct, both models exhibiting very strong zero-shot performance and a downward trend especially with 3 and 5 examples.
@@ 6.4 RoCulturaBench
Results on RoCulturaBench are presented in Table 7. Taking an English foundational model and adapting it to a new language by performing instruction fine-tuning is not enough to properly understand the nuances and realities of a culture. Testament stands the competitive performance of RoLlama2: starting from an arguably weaker model but also pre-training it performs on par with newer and stronger models that have only been fine-tuned.
We argue that, as expected, cultural information and social knowledge are primarily injected into the model in the pre-training phase, not in the fine-tuning phase, especially when fine-tuning exploits non-natively data (i.e., written in English and then translated). As in the case of MT-Bench, Llama3 outperforms other models by a noteworthy margin, with only GPT-3.5 being stronger. For more detailed results and comparisons, see Appendix Section A.2. In addition, results obtained using the LLM-as-a-judge approach are not the be-all and end-all of evaluations, especially regarding non-English language and culture. We leave human evaluation and other approaches (e.g., training a language-specific judge) for future work.
Model RoMT-Bench RoCulturaBench
RoLlama2-7b-Instruct 4.43 4.08
RoLlama2-7b-Instruct-DPO 4.61 4.80
RoMistral-7b-Instruct 5.29 3.99
RoMistral-7b-Instruct-DPO 5.88 4.72
RoLlama3-8b-Instruct 5.38 3.81
RoLlama3-8b-Instruct-DPO 5.87 4.40
RoLlama3.1-8b-Instruct 5.42 3.55
RoLlama3.1-8b-Instruct-DPO 6.21 4.42
RoGemma-7b-Instruct 5.24 3.51
RoGemma-7b-Instruct-DPO 5.48 3.94
Table 8: Performance of aligned RoLLMs. Bold denotes the best results for each benchmark.
6.5 Human Alignment
After manually inspecting part of the results for RoMTBench and RoCulturaBench, we found that our models tend to have a more rigid answer. The other models have a more conversational tone, which is more in line with the reference answers and general human preferences. This sometimes rigid tone of RoLLMs stems mostly from the instruction dataset used for fine-tuning and the lack of human preference alignment mechanisms. To test this, we decided to translate the HelpSteer dataset Wang et al. (2024) containing human preference data and use it to perform Direct Preference Optimization Rafailov et al. (2024) (DPO) on our models.
Results are presented in Table 8. Using only 15k samples, we notice a significant increase in performance across the board, proving the essential role of human alignment for LLMs. We leave a more thorough approach for human alignment, including using more data, to future work.